A farmer mixes a 20% fertilizer solution with a 40% fertilizer solution to make 8 quarts of a 33.75% fertilizer solution. How many quarts of each does she use?
Answer by lwsshak3(11628)

A farmer mixes a 20% fertilizer solution with a 40% fertilizer solution to make 8 quarts of a 33.75% fertilizer solution. How many quarts of each does she use?
**
let x=quarts of 20% solution
8-x=quarts of 40% solution
..
20%x+40%(8-x)=33.75%*8
.2x+3.2-.4x=2.7
.2x=.5
x=2.5
8-x=5.5
ans:
quarts of 20% solution used=2.5
quarts of 40% solution used=5.5